Sometimes we need to find out how long it will take money to grow to some particular amount. For example, if a company's sales are growing at 10% per year, how long will it takes sales to triple? This answer is 11.53

***The answer I need help with is as followers. If you want the investment to double in 8 years, what interest rate must it earn?
Answer #1

Amount Invested = $1
Desired Sum = $2
Period = 8 years

Let Interest Rate be i%

Amount Invested * (1 + Interest Rate)^ Period = Desired Sum
$1 * (1 + i)^8 = $2
(1 + i)^8 = 2
1 + i = 1.0905
i = 0.0905 or 9.05%

So, you must earn an interest rate of 9.05%